P(4,9) ,Q(3,2),R(13,-3) and S(11,13) are four points in a plane and L is the mid point of RS . the line through Q perpendicular to RS meets PR at M find :
1 the equation of the lines QL and PR
2 the coordinates of the point of intersection N ,of QL and PR
3 the length of QM
4 the area of triangle PQN
'
1
Expert's answer
2019-09-03T11:02:30-0400

1. Find the coordinates of L:


"x=\frac{13+11}{2}=\frac{24}{2}=12""y=\frac{-3+13}{2}=\frac{10}{2}=5"


Find the slope of the line QL:


"m_{QL}=\frac{y_2-y_1}{x_2-x_1}=\frac{5-2}{12-3}=\frac{3}{9}=1/3"

Substitute the slope and point Q in the slope intercept form of the equation ("y=mx+b" ) and solve for b:


"2=\frac{1}{3}\cdot3+b""2=1+b""b=1"

The equation of the line QL is: "y=\frac{1}{3}x+1".


Find the slope of the line PR:


"m_{PR}=\frac{-3-9}{13-4}=\frac{-12}{9}=-\frac{4}{3}"

Substitute the slope and point P in the slope intercept form of the equation and solve for b:


"9=-\frac{4}{3}\cdot 4+b""9=-\frac{16}{3}+b""b=\frac{43}{3}"

The equation of the line PR is: "y=-\frac{4}{3}x+\frac{43}{3}".


2. If N is the point of intersection of QL and PR, then its coordinates satisfy the system of equations:


"\begin{cases} y=\frac{1}{3}x+1 \\ y=-\frac{4}{3}x+\frac{43}{3} \end{cases}""\frac{1}{3}x+1=-\frac{4}{3}x+\frac{43}{3}""\frac{5}{3}x=\frac{40}{3}""x=\frac{40}{3}:\frac{5}{3}=8""y=\frac{8}{3}+1=\frac{11}{3}"

Hence "N(8,\frac{11}{3})" - the point of intersection of QL and PR.


3. Find the slope of the line RS:


"m_{RS}=\frac{13-(-3)}{11-13}=\frac{16}{-2}=-8"

Since the line MQ is perpendicular to the line RS,  their slopes have a particular relationship to each other:


"m_{MQ}=-\frac{1}{m_{RS}}=\frac{1}{8}"

Find the equation of the line MQ. Substitute the slope and point Q in the slope intercept form of the equation and solve for b:

"2=\frac{1}{8}\cdot{3}+b""b=2-\frac{3}{8}""b=\frac{13}{8}"

The equation of the line MQ is: "y=\frac{1}{8}x+\frac{13}{8}"

If N is the point of intersection of MQ and PR, then its coordinates satisfy the system of equations:


"\begin{cases} y=\frac{1}{8}x+\frac{13}{8} \\ y=-\frac{4}{3}x+\frac{43}{3} \end{cases}"

"\begin{cases} 8y=x+13 \\ 3y=-4x+43 \end{cases}""\begin{cases}x=8y-13 \\ 3y=-4x+43 \end{cases}"

"3y=-4(8x-13)+43"

"3y=-32x+52+43"

"35y=95"

"y=\frac{19}{7}"

"x=8\cdot\frac{19}{7}-13=\frac{61}{7}"

The coordinates of M are "(\frac{61}{7},\frac{19}{7})"

Then the length of QM:


"QM=\sqrt{(\frac{61}{7}-3)^2+(\frac{19}{7}-2)^2}"


"QM=\sqrt{(\frac{40}{7})^2+(\frac{5}{7})^2}=\sqrt{\frac{1625}{49}}=\frac{5\sqrt{65}}{7}"


4. Find the area of the triangle using the formula:


"A=\lvert \frac{x_1(y_2-y_3)+x_2(y_3-y_1)+x_3(y_1-y_2)}{2}\rvert"

"A=\lvert\frac{4(2-\frac{11}{3})+3(\frac{11}{3}-9)+8(9-2)}{2}\rvert"

"A=\lvert\frac{-\frac{20}{3}-16+56}{2}\rvert=\frac{50}{3}"